OpenStudy (anonymous):

simplify by factoring square root 24

OpenStudy (anonymous):

type an exact answer using radical as needed

OpenStudy (anonymous):

24=4*6 so sqrt(4*6)=(sqrt(4))*(sqrt(6))=2sqrt(6)

OpenStudy (anonymous):

so the exact answer is 2sqrt 6

OpenStudy (anonymous):

yea... unless you can use decimals

OpenStudy (usukidoll):